(b) Find the equation of the sphere, which contains the circle
x^2 + y^2 + z^2 = 9, 3x + 3y + 3z = 5 and
passes through the origin.

Answer by venugopalramana(3286)

(b) Find the equation of the sphere, which contains the circle
x^2 + y^2 + z^2 = 9, 3x + 3y + 3z = 5 and passes through the origin.
THAT IS CONTAINS CIRCLE
x^2 + y^2 + z^2 - 9=0=P SAY ,AND 3x + 3y + 3z - 5 =0=L SAY
ANY SPHERE THROUGH THE ABOVE CIRCLE IS GIVEN BY
P+KL=0...WHERE K IS A CONSTANT TO BE FOUND..WE ARE GIVEN IT PASSES THROUGH ORIGIN.HENCE
0^2+0^2+0^2-9+K(0+0+0-5)=0
-9-5K=0...........5K=-9
K=-9/5
HENCE EQN OF SPHERE IS
5(X^2+Y^2+Z^2-9)-9(3X+3Y+3Z-5)=0
5X^2+5Y^2+5Z^2-45-27X-27Y-27Z+45=0
5X^2+5Y^2+5Z^2-27X-27Y-27Z=0